# Normalizing Flows for Gaussian Process Regression under Hierarchical Shrinkage Priors

### v1

@@ 4 Simulation Study
The results are presented in Figure 4. Generally, the proposed normalizing flow approach outperforms both the ML and mean field method, especially at higher sparsity levels and dimensions, although all methods perform comparatively when the ratio of data points to covariates is high. Even when performing equally well in terms of the median LPDS, the ML and mean field approach often displays higher variability, potentially indicating that it is less robust in estimating predictive densities. The triple gamma prior, with its aggressive shrinkage and heavy tails, performs particularly well in high sparsity settings, while still being competitive in lower sparsity settings. It is important to note that the cost for this increased predictive performance is a higher computational burden. While the normalizing flow approach is computationally efficient for a fully Bayesian approach, it is still remains significantly slower than the ML approach.
@@ 5 Concluding Remarks and Outlook
This work has contributed to the literature on Gaussian Process Regression (GPR) in two key ways. First, hierarchical shrinkage in the form of the hierarchical triple gamma prior was investigated as a mechanism to mitigate the curse of dimensionality in high-dimensional regression settings. The prior demonstrated the ability to effectively exclude irrelevant covariates while maintaining flexibility in model size. Second, a novel approach to estimating GPR models was introduced, leveraging normalizing flows within a variational inference framework. This approach enables the approximation of complex posterior distributions in a computationally efficient manner, capitalizing on advances in GPU computing. The proposed methodology outperformed a standard maximum likelihood-based approach in simulation studies, particularly in settings characterized by high dimensionality and sparsity.
The use of normalizing flows for posterior approximation remains a relatively novel concept in the context of GPR models, and there are several promising avenues for future research. First, the choice of transformation is critical to the method’s performance, and exploring alternatives such as the Inverse Autoregressive Flow (Kingma et al., 2016) could yield further improvements. Recent studies, including Kong and Chaudhuri (2020), suggest that Sylvester flows may lack the expressivity of other options, underscoring the potential benefits of investigating more flexible transformations. Second, while the methodolgy is currently able to scale to datasets of reasonable size, extending it to sparse GPR models (see, e.g., Quinonero-Candela and Rasmussen, 2005, for an overview) could enable its application to truly large datasets. However, such an extension would need to address the computational challenges associated with estimating a large number of inducing points, which would significantly increase the dimension of the posterior, thereby also increasing the size of the approximating normalizing flow. Finally, extending the approach to non-Gaussian outcomes may be of interest. The absence of a closed-form marginalized likelihood in this setting would necessitate inference of the latent function $f(\mathbf{x})$, increasing posterior dimensionality and potentially reducing computational efficiency. These challenges notwithstanding, the proposed methodology offers a promising foundation for future research in GPR estimation.
{funding}
This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/J4806.

### v2

@@ 4 Simulation Study
The results are presented in Figure 5 for $\rho=0.5$ and Figure 6 for $\rho=0.9$. Overall, the NF approach seems to have the largest advantage when there are many covariates as well as a high degree of sparsity, as can be seen in the second and third row of the two figures. In such cases, the combination of aggressive shrinkage and rich posterior approximation of the normalizing flows seems to provide the largest benefit. While the prior is the same in the mean-field approach, the rougher mean-field approximation can not take advantage of this benefit to the same degree. Maximum likelihood performs well when the ratio of covariates to observations is not too unfavorable, but falls apart when the covariate space becomes too high-dimensional.
Zooming in on the results, the first rows show that all methods perform similarly when there are only 10 covariates, with only slight dips in performance from the mean-field and maximum likelihood approach when the sparsity is 0.5, $N=100$ and $d=10$. As soon as more observations are available, all methods perform approximately the same. The results are similar in the second row, albeit starker: when the amount of data points are low and sparsity is high, the NF approach can handily outperform the other approaches, see the $N=100,d=50$ cases. When multicollinearity is high ($\rho=0.9$) and selecting incorrect covariates that correlate strongly with the truly relevant ones can lead to similar predictive performance, this advantage becomes smaller. Under lower multicollinearity ($\rho=0.5$), the NF approach is often far better than competing approaches, as correct selection of covariates is more important.
When the problem becomes more high-dimensional in the third row ($d=200$), the results are similar again, with some caveats. In the lower multicollinearity setting, all methods perform poorly, leading to roughly equal predictive performance. This is the result of all methods predicting the marginal distribution of the data for new data points, as there is too much noise to effectively filter out signal. The only exception lies in the $N=2500,d=200$ scenario with sparsity = 0.9, where the aggressive shrinkage of the triple gamma with $a=c=0.1$ can effectively filter out irrelevant covariates and thereby give better predictions than competing methods. In the higher multicollinearity setting, the NF approach is able to consistently outperform the competing methods, with this advantage becoming larger the sparser the true underlying covariates are.
The results seem to imply that the combination of strong shrinkage in the kernel as well as the possibility to approximate complex posteriors is most advantageous when the conditions are adverse, namely in situations that are high-dimensional, highly sparse and with covariates that are strongly correlated with one another. As all methods perform similarly to one another when conditions are more favorable, there seems to be little disadvantage in using the proposed approach over others, making it a fairly "safe" choice for estimating GPR models.
5 Application to BACE-1 Inhibitor Data
Figure 7: Log Predictive Density Scores (LPDS) across 10-fold cross-validation for Maximum Likelihood (ML), mean-field (MF) approximation under a horseshoe (MF: HS) and triple gamma prior (MF: TG) and the proposed approach under a horseshoe prior (NF: HS) and triple gamma prior (NF: TG) for the BACE-1 inhibitor dataset
Figure 8: Posterior distribution of the model parameters $\theta_{1},\dots,\theta_{d}$ for the BACE-1 inhibitor dataset under triple gamma prior with $a=c=0.1$. Sorted in descending order of posterior median.
$\beta$-secretase (BACE-1) is widely recognized as a key therapeutic target for Alzheimer’s disease (vassar2011beta), as the inhibition of BACE-1 has been shown to reduce the formation of amyloid plaques in the brain (rossner2001neuronal), which are associated with the disease. The BACE-1 inhibitor dataset (wu2018moleculenet) contains quantitative binding results for 1513 potential inhibitors of BACE-1, described by 589 molecular descriptors. The target is to predict the $\text{pIC}_{50}$ value, which is the negative logarithm of $\text{IC}_{50}$, which in turn is the concentration of an inhibitory substance (e.g., a drug) required to reduce a specific biological process or component’s activity by 50% in vitro.
The dataset was preprocessed by removing all covariates with a variance of zero, scaling all non-binary covariates to have a mean of zero and a standard deviation of one and centering the response variable. This left 363 explanatory variables. All models were evaluated by their out-of-sample LPDS based on 10-fold cross-validation, with each training set containing 90% of the data and the corresponding test set containing 10%. The models compared are the same as those used in the simulation study, with the only difference being the amount of iterations allowed for convergence for both variational inference approaches. The normalizing flow models were run for 5000 iterations to ensure convergence, although this was typically achieved earlier. The mean-field approach was run for 10000 iterations, showing worse convergence than the normalizing flow approach.
The results are presented in Figure 7. The proposed normalizing flow approach outperforms both the maximum likelihood and mean-field method, especially when using the more aggressive triple gamma prior. The mean-field approach often did not converge to a useful result, running the maximum number of iterations with the ELBO stagnating or even diverging. To gain some insight into the results generated by the normalizing flow approach, the posterior distribution of $\theta_{1},\dots,\theta_{d}$ under the triple gamma prior with $a=c=0.1$ are visualized in Figure 8, sorted in descending order of posterior median. Many of the $\theta_{j}$ have a posterior that concentrates around zero, indicating that the corresponding covariates do not meaningfully contribute to the model. This is in line with the expectation that many molecular descriptors in the dataset are not directly related to BACE-1 inhibition. On the flip side, the parameters that are further away from zero often display some uncertainty and, when looking at pair plots, dependencies between them can be observed. Such intricacies could not be accurately captured by either competing method and most likely contributed to the better predictive performance of the proposed approach.
@@ 6 Concluding Remarks and Outlook
This work has contributed to the literature on GPR in two key ways. First, hierarchical shrinkage in the form of the hierarchical triple gamma prior was investigated as a mechanism to mitigate the curse of dimensionality in high-dimensional GPR settings. The prior demonstrated the ability to effectively exclude irrelevant covariates while maintaining flexibility in model size. Second, a novel approach to estimating GPR models was introduced, leveraging normalizing flows within a variational inference framework. This approach enables the approximation of complex posterior distributions in a computationally efficient manner, capitalizing on advances in GPU computing. The proposed methodology outperformed a standard maximum likelihood-based approach as well as a mean-field based approach in simulation studies, particularly in settings characterized by high dimensionality and sparsity. This advantage was also borne out in an application to BACE-1 inhibitor data, which is charaterized by a large amount of covariates and high sparsity. Here, the combination of normalizing flows and aggressive shrinkage was instrumental in providing increased predictive performance. Additionally, the methods presented here are implemented in the R package shrinkGPR (shrinkGPR), which is publicly available on CRAN.
The use of normalizing flows for posterior approximation remains a relatively novel concept in the context of GPR models, and there are several promising avenues for future research. First, the choice of transformation is critical to the method’s performance, and exploring alternatives such as the Inverse Autoregressive Flow (kingma2016improved) could yield further improvements. Recent studies, including kong2020expressive, suggest that Sylvester flows may lack the expressivity of other options, underscoring the potential benefits of investigating more flexible transformations. Second, while the methodolgy is currently able to scale to datasets of reasonable size, extending it to sparse GPR models (see, e.g., quinonero2005unifying, for an overview) could enable its application to truly large datasets. However, such an extension would need to address the computational challenges associated with estimating a large number of inducing points, which would significantly increase the dimension of the posterior, thereby also increasing the size of the approximating normalizing flow. Finally, extending the approach to non-Gaussian outcomes may be of interest. The absence of a closed-form marginalized likelihood in this setting would necessitate inference of the latent function $f(\mathbf{x})$, increasing posterior dimensionality and potentially reducing computational efficiency. These challenges notwithstanding, the proposed methodology offers a promising foundation for future research in GPR estimation.
7 Disclosure statement
The author declares that they have no competing interests.
8 Data Availability Statement
The BACE-1 data is, as of 2025, publicly available at https://moleculenet.org/datasets-1.
SUPPLEMENTARY MATERIAL
R-package shrinkGPR:
R-package shrinkGPR, which implements the methods described in the article, is available on CRAN at https://CRAN.R-project.org/package=shrinkGPR.